# Open Datasheets: Machine-readable Documentation for Open Datasets and Responsible AI Assessments

### v1

@@ Abstract
This paper introduces a no-code, machine-readable documentation framework for open datasets, with a focus on Responsible AI (RAI) considerations. The framework aims to improve the accessibility, comprehensibility, and usability of open datasets, facilitating easier discovery and use, better understanding of content and context, and evaluation of dataset quality and accuracy. The proposed framework is designed to streamline the evaluation of datasets, helping researchers, data scientists, and other open data users quickly identify datasets that meet their needs and/or organizational policies or regulations. The paper also discusses the implementation of the framework and provides recommendations to maximize its potential. The framework is expected to enhance the quality and reliability of data used in research and decision-making, fostering the development of more responsible and trustworthy AI systems.
1 Introduction
@@
Machine-readable documentation, which refers to documentation that is structured in a format easily readable and processed by computers, plays a vital role in improving the accessibility, comprehensibility, and usability of open datasets. It has the potential to enable easier discovery and use of datasets, facilitate a better understanding of the content and context of a dataset, simplify the integration of datasets from various sources, and help with the evaluation of the quality and accuracy of datasets heger2022understanding . With the exponential growth of open data, machine-readable metadata is becoming increasingly important for making this data easily discoverable and useable contaxis2022ten .
@@
As open data is progressively employed in AI applications mckinsey_opendata ; BRINKHAUS2023102542 , it is imperative for researchers to consider the potential implications of these open datasets from a Responsible AI (RAI) perspective heger2022understanding . It is widely recognized that AI systems can inherit biases and other flaws from the data they are trained on  heger2022understanding ; buolamwini2018gender , and with AI continuing to permeate various aspects of our daily lives, the transparency, reliability, trustworthiness, and fairness of AI systems becomes paramount. This begins with the quality of the data they are trained on paullada2021data . Understanding the data being used in an AI system is crucial to uphold responsible AI implementations, including its source, the methodologies used in its collection, and other relevant factors discussed in this paper.
@@
Evaluating a dataset involves multiple distinct goals or aspects. One is to determine if a dataset complies with organizational policies and/or regulations, such as ensuring appropriate consent was obtained. The other is to ascertain if a dataset is suitable for the task at hand, such as whether the labels are valid proxies for the real-world phenomena a model is being designed to predict. While researchers can often evaluate an open dataset and make determinations about some of its characteristics, such as demographic distribution, as part of their exploratory data analysis (EDA), there are properties that cannot be learned through data exploration, such as exactly how demographic information was obtained from data subjects or whether and how their consent was obtained. This makes searching for and evaluating datasets an overwhelming, if not impossible, task. It can lead to investing a significant amount of time in evaluating a dataset only to realize later that it was not collected properly and/or does not have a proper representation from an RAI perspective to meet their organizational policies, among other considerations. Such mistakes can be costly, with potential harm to individuals, and significant losses for organizations iapp_webscrape_data ; price2019privacy ; trivedi2019risks ; forbes_cost_bad_data ; entrepreneurs_cost_bad_data ; hbs_data_integrity_importance ; gartner_cost_bad_data .
@@
2 Related Work
Our proposed framework, Open Datasheets, builds on the line of work started in “Datasheets for Datasets” gebru2021datasheets , and followed by “Understanding machine learning practitioners’ data documentation perceptions, needs, challenges, and desiderata” heger2022understanding . These papers emphasized the importance of providing comprehensive documentation for datasets to ensure transparency, reproducibility, and accountability in the use of data.
Expanding upon these ideas, the Open Datasheets framework introduces a machine-readable metadata format for open datasets. This format provides detailed information about the datasets, including responsible AI considerations, and adheres to the 7 design principles outlined in the research conducted by Heger et al. heger2022understanding , as discussed in the design (section 3) and implementation (section 4) sections.
To contribute to the existing field of data documentation, the framework takes into account other published frameworks. The aim is not to reinvent the wheel but rather to provide an applied and efficient open framework that is user-friendly for non-developers and other open data publishers and users. Additionally, the framework aims to integrate with existing open platforms and support conversion to other standard formats, such as JSON-LD.
Among the evaluated frameworks are Datacards pushkarna2022data , which align with most of the principles of the “Datasheets for Datasets” paper but lack machine-readable outputs, integration with existing open platforms, and automation features. Dataset Nutrition Labels holland2020dataset ; chmielinski2022dataset also align with these principles but suffer from similar limitations, including a closed system and mandatory registration. Another specialized framework called Data statements bender2018data for natural language processing (NLP) focuses specifically on NLP datasets.
@@ 2 Related Work
Other documentation frameworks and formats, such as those utilized by HuggingFace mcmillan2021reusable ; lhoest2021datasets , were also considered but not included here due to their specialization and association with proprietary platforms. The evaluation focused on flexible published frameworks that can be adapted, are open, and align closely with the 7 principles proposed by Heger et al. heger2022understanding .
3 Design
@@
The Open Datasheets framework is designed to streamline the documentation of open datasets. It aims to foster the inclusion of crucial information that assists users in comprehending potential biases, privacy concerns, and other elements of responsible AI. This is achieved by striking a balance between automation and thoughtful deliberation in the documentation process  heger2022understanding . The ultimate objective is to enhance data documentation for open datasets, improve their discoverability, and promote their reliability, fairness, and transparency when used for AI models.
@@
The Open Datasheets format specification is rooted in the Datapackage standard, a product of the Frictionless Data project FrictionlessData2023 . This open-source initiative offers tools and standards for data management. The Datapackage specification, accessible on GitHub frictionlessdata_spec_github , is a user-friendly JSON based format that acts as a container for describing a set of data. Also, it supports different data types, including complex ones such as tabular and geographic data, allowing users to document the basic information of a dataset in a way that is tailored to their specific needs. This standard has been widely adopted on open data platforms such as  https://data.world.
@@
The Open Datasheets framework utilizes the Datapackage format specifications to consistently document basic dataset details. This approach enables users to easily comprehend the information on the datasheets by avoiding specialized terminology and adhering to the standard language and usage of the JSON format. This format is machine-readable and seamlessly integrates into data workflows, thereby enhancing data discovery, access, and usage. The adoption of this specification’s fosters compatibility and uniformity in the realm of data documentation.
@@
  "name" : "a-unique-human-readable-and-url-usable-identifier",
  "title" : "A nice title",
@@
Although the Datapackage specification is beneficial for general data documentation, it may not suffice for organizations that need to comply with regulatory standards and ensure responsible AI practices. To cater to this need, the Open Datasheets framework incorporates principles from “Datasheets for Datasets” gebru2021datasheets . These principles, such as detailed descriptions of a dataset’s privacy implications, are integrated into a machine-readable format. This additional information is crucial for organizations to make informed decisions about dataset usage and ensure compliance with their policies.
Moreover, the Open Datasheets Framework adheres to the 7 design principles outlined in section 5.2 of Heger et al research heger2022understanding . It explicitly connects data documentation with responsible AI through inline guidance based in part on Microsoft’s Aether data documentation template aether_doc . The framework is practical, integrating with GitHub and providing a user-friendly interface. It adapts to different contexts, allowing customization. Automation is balanced with responsibility, automating foundational metadata extraction while guiding users on responsible AI metadata. It clarifies the target audience to data publishers, focusing on potential users of open datasets. Standardization is achieved through a machine-readable format, and integration with GitHub promotes collaboration. The framework seamlessly integrates into existing tools and workflows, associating documentation with the data publication and development lifecycle on GitHub.
@@
4 Implementation
To achieve a no-code solution, the framework implements a user-friendly web application on GitHub Pages 1, a managed service by the GitHub Platform to host static websites on GitHub repositories. This implementation ensures the longevity and support of the framework. The web application features a wizard-style interface that assists in standardizing the documentation according to the framework’s metadata format. It includes metadata parsers for common data file formats on the GitHub platform roman2023open , such as CSV, TSV, JSON, and others. These parsers extract metadata related to the file structure, including field names, types, and sample values. General metadata about the data file, such as filename, encoding, and size, is also extracted. This approach allows data publishers to focus solely on filling in the responsible AI metadata.
@@ 4 Implementation
The integration with the GitHub Platform, one of the largest open data platforms roman2023open , promotes openness and community. Currently, the web application enables seamless documentation of datasets on GitHub, enhancing discoverability and transparency for datasets hosted on the platform. However, the Open Datasheets framework is not limited to documenting datasets exclusively on GitHub. It provides the flexibility to create documentation for datasets and download the metadata file for publishing on any preferred platform. Moreover, the framework is openly available on GitHub, ensuring easy accessibility for data publishers.
The Open Datasheets framework simplifies the data documentation process for data publishers and promotes the inclusion of responsible AI elements. By automating the extraction of foundational metadata from data files and providing visual guidance for responsible AI considerations, the framework reduces the time and effort required for extensive documentation. This addresses the reluctance of data publishers to write lengthy documentation, as highlighted in Heger et al research  heger2022understanding .
For data users, the Open Datasheets framework offers detailed and machine-readable documentation, enabling informed decisions regarding dataset selection and use. The framework encourages data publishers to document potential biases and privacy implications associated with the dataset. This documentation empowers data users to evaluate the reliability and fairness of the dataset. The comprehensive documentation provided by the framework enhances transparency and trustworthiness, allowing data users to assess the dataset’s quality, understand its limitations, and ensure it aligns with their ethical standards and requirements.
@@
The Open Datasheets framework metadata format comprises two main sections: the dataset foundational metadata and the responsible AI metadata. The dataset foundational metadata extends the Datapackage standard and includes essential information such as the dataset’s name, title, licenses, sources, and a comprehensive list of the data resources within the package. On the other hand, the responsible AI metadata follows the principles outlined in the "Datasheets for datasets" research paper. It provides information about the data’s origin, processing methods, privacy implications, potential biases, and other relevant aspects. These components are further discussed in the following sections, specifically Section  4.1 and Section  4.2.
4.1 Datapackage
@@
4.2 Responsible AI
The Responsible AI module of the Open Datasheets framework focuses on documenting aspects that can impact user privacy, security, and biases that can undermine the trustworthiness of AI systems when used for training AI models. Evaluating datasets is crucial for creating responsible AI systems because the quality of the data directly affects the accuracy and fairness of the resulting system. Biases in the data can lead to biased outcomes, which can have serious consequences for individuals and society obermeyer2019dissecting ; chen2023human ; mckinsey_bias . For instance, if an AI system used for hiring decisions is trained on biased data, it may perpetuate discrimination against certain groups.
@@ 4.2 Responsible AI
However, it is important to note that human review may still be necessary to make decisions based on organizational policies. The Open Datasheets framework is not a substitute for human judgment but rather a tool to assist in the evaluation of responsible AI. By combining human expertise with machine-readable metadata, organizations can create more responsible and trustworthy AI systems.
With this objective in mind, the framework encompasses the following areas:
  • Privacy ( 4.2.1)
  • Data Access ( 4.2.2)
  • Collection Procedures ( 4.2.3)
  • Processing Procedures ( 4.2.4)
  • Use Cases ( 4.2.6)
4.2.1 Privacy and Use Terms
@@
is a key aspect of privacy assessments. It refers to the protection of personal information from unauthorized access, use, or disclosure. Therefore, having this incorporated into the framework helps the users of the dataset understand the measures that were taken, if any, by the publishers to avoid the disclosure of personal information in their dataset.
@@
in the context of privacy, refers to the level of sensitivity or potential harm associated with certain types of personal information included in a dataset. This includes sensitive attributes such as race, sex, religion, sexual orientation, health information, financial data, and other personally identifiable information (PII). These sensitive attributes have the potential to cause harm or discrimination if mishandled, accessed, or disclosed without proper consent or safeguards.
@@
4.2.2 Data Access
Including information on how to access a dataset it is also essential for privacy and security considerations. One key reason is to clearly communicate whether the data can be accessed anonymously or if it requires registration. This information helps potential users understand the accessibility of the data and determine if they meet the necessary criteria to access it.
@@ 4.2.2 Data Access
Another important reason for specifying data access methods in the dataset documentation is to establish clear terms of use. Explicitly accepting and agreeing to use terms is essential to protect the dataset and maintain data security. This helps prevent unauthorized access or misuse of the data, preserving its integrity and privacy.
In line with these considerations, the Open Datasheets framework incorporates a module that enables data publishers to clearly specify how people can access the data. This module helps users understand the accessibility of the data and determine if it suits their needs. Additionally, it enables dataset providers to establish terms of use that protect the data and maintain data security. By providing this information, data providers can promote transparency, accountability, and responsible data usage.
4.2.3 Collection Procedures
@@
Additionally, documenting data collection procedures enables the replication of the process in the future. This is crucial as it allows researchers to verify the results of an AI system trained on that data and confirm that the system is functioning as intended. Furthermore, documenting data collection helps data users ensure that the process adheres to ethical and legal standards.
The framework takes into consideration the methods, consent forms utilized for data subjects and the contributors of the collection procedures.
@@ Methods
describe who perform the data collection, if an individual, an organization or a third-party. This helps data users understand if the data publisher was not the entity that collected the data.
4.2.4 Processing Procedures
@@
4.2.5 Update Procedures
Datasets can be categorized as either static or periodically updated. A static dataset remains unchanged over time, while a periodically updated dataset undergoes regular updates to incorporate new data or revisions. When choosing between these options, researchers consider their specific objectives. If the research requires analyzing historical data or examining trends over a specific period, a static dataset may be suitable. Conversely, if the research involves analyzing real-time or evolving phenomena, a periodically updated dataset would be more appropriate. By documenting the procedures for updating the dataset, data users can gain insights into the timing and nature of the updates. This documentation promotes transparency and builds trust in the data, as users can verify its currency and reliability.
@@ 4.2.5 Update Procedures
describes the approach for updating the dataset, such as incrementally or through other methods.
@@ 4.2.6 Use Cases
Overall, documenting the permissible and impermissible uses of a dataset is a crucial step in ensuring responsible AI systems. It provides guidelines for utilizing the data in a responsible and ethical manner, while also holding individuals accountable for any issues that may arise with the AI.
@@
4.3 Evaluation
To assess the design and implementation of the Open Datasheets framework, we collaborated with a diverse group of open dataset producers, including data scientists, researchers, analysts, and program managers from Microsoft. Our objective was to gain valuable insights into their publishing processes and the level of documentation they employed. Moreover, we aimed to explore how perspectives varied across different roles, such as analysts versus data scientists.
We conducted an initial evaluation during the early stages of the Open Datasheets web application. This evaluation involved leveraging the framework for three real open data publications clinical_sum_github ; clandestino_github ; rtp_lx_github on GitHub from producers with diverse backgrounds, including non-developer researchers and data scientists. While providing a detailed report of this study is beyond the scope of this paper, we discovered that participants encountered challenges in comprehending the value of RAI documentation and determining the required information. These difficulties were exacerbated by the lack of inline guidance within the tool, which aligns with the conclusions of Heger et al.’s research heger2022understanding . Furthermore, automation features were not available during the evaluation period. Participants primarily focused on including only essential information for datasets that consisted of multiple data resources.
Based on these insights, we improved the design and implementation to better balance automation and inline guidance, as outlined in the design and implementation sections of this paper.
5 Conclusion

### v2

@@ Abstract
This paper introduces a no-code, machine-readable documentation framework for open datasets, with a focus on responsible AI (RAI) considerations. The framework aims to improve comprehensibility, and usability of open datasets, facilitating easier discovery and use, better understanding of content and context, and evaluation of dataset quality and accuracy. The proposed framework is designed to streamline the evaluation of datasets, helping researchers, data scientists, and other open data users quickly identify datasets that meet their needs and organizational policies or regulations. The paper also discusses the implementation of the framework and provides recommendations to maximize its potential. The framework is expected to enhance the quality and reliability of data used in research and decision-making, fostering the development of more responsible and trustworthy AI systems.
Introduction
@@
Machine-readable documentation, which refers to documentation that is structured in a format easily readable and processed by computers, plays a vital role in improving the accessibility, comprehensibility, and usability of open datasets. It has the potential to enable easier discovery and use of datasets, facilitate a better understanding of the content and context of a dataset, simplify the integration of datasets from various sources, and help with the evaluation of the quality and accuracy of datasets[1]. With the exponential growth of open data, machine-readable metadata is becoming increasingly important for making this data easily discoverable and useable[2].
@@
As open data is progressively employed in AI applications[3, 4], it is imperative for researchers to consider the potential implications of these open datasets from a responsible AI (RAI) perspective[1]. It is widely recognized that AI systems can inherit biases and other flaws from the data they are trained on [1, 5], and with AI continuing to permeate various aspects of our daily lives, the transparency, reliability, trustworthiness, and fairness of AI systems becomes paramount. This begins with the quality of the data they are trained on[6]. Understanding the data being used in an AI system is crucial to uphold responsible AI implementations, including its source, the methodologies used in its collection, and other relevant factors discussed in this paper.
@@
Evaluating a dataset involves multiple distinct goals or aspects. One is to determine if a dataset complies with organizational policies and/or regulations, such as ensuring appropriate consent was obtained. The other is to ascertain if a dataset is suitable for the task at hand, such as whether the labels are valid proxies for the real-world phenomena a model is being designed to predict. While researchers can often evaluate an open dataset and make determinations about some of its characteristics, such as demographic distribution, as part of their exploratory data analysis (EDA), there are properties that cannot be learned through data exploration, such as exactly how demographic information was obtained from data subjects or whether and how their consent was obtained. This makes searching for and evaluating datasets an overwhelming, if not impossible, task. It can lead to investing a significant amount of time in evaluating a dataset only to realize later that it was not collected properly and/or does not have a proper representation from an RAI perspective to meet their organizational policies, among other considerations. Such mistakes can be costly, with potential harm to individuals, and significant losses for organizations[7, 8, 9, 10, 11, 12, 13].
@@
Related Work
In recent years there have been calls for comprehensive documentation of the datasets used to train and evaluate AI systems[14, 1, 15, 16, 17, 18]. Several data documentation frameworks and tools have been proposed with the goal of encouraging thoughtful reflection on datasets and transparency about their makeup and creation process. Datasheets for datasets [14] is a documentation framework designed to encourage dataset creators to reflect on the implicit and explicit choices behind their data and to enable those interested in using a dataset to train or evaluate their system to make more informed decisions about its appropriateness. Developed concurrently, data nutrition labels [15, 19] also include information about a dataset’s provenance and makeup. The authors created a publicly available tool to create a nutrition label that highlights usage restrictions and potential harms that may arise from the dataset’s use. Data statements for natural language processing [16] are designed specifically for natural language datasets and include specialized information such as speaker demographics and language variety. More recently, researchers at Google released the Data Cards Playbook[20], while researchers at Microsoft released the Aether Data Documentation Framework, a variant of datasheets for datasets adapted to meet the needs of industry practitioners[1].
While research shows that well-designed data documentation can be effective at helping to identify ethical issues[21], good documentation can be challenging to create. In a study of industry practitioners’ data documentation perceptions, needs, challenges, and desiderata, Heger et al. [1] found that practitioners creating dataset documentation had trouble making connections between the questions they were asked to answer and their RAI implications and difficulty providing information that someone unfamiliar with their datasets would need to understand the data. Based on their findings, they derived seven design requirements that data documentation frameworks should satisfy. Briefly, these include 1) making the connection to RAI more explicit; 2) making data documentation frameworks practical; 3) adapting data documentation frameworks to different contexts; 4) supporting simple tasks with automation without automating away responsibility; 5) clarifying the target audience for the documentation; 6) standardizing and centralizing data documentation; and 7) integrating data documentation frameworks into existing tools and workflows.
Inspired by and building on this line of work, our proposed framework, Open Datasheets, introduces a machine-readable metadata format for open datasets. Drawing on questions included in datasheets for datasets [14] and the Aether data documentation template [22], it includes detailed information about the datasets, including responsible AI considerations. Furthermore, it adheres to the seven design principles of Heger et al.[1], as discussed in the design and implementation sections. Unlike existing tools, the framework aims to integrate with existing open platforms and support conversion to other standard formats, such as JSON-LD, providing an applied and efficient open framework that is user-friendly for non-developers and other open data publishers and users.
@@
Other documentation frameworks and formats, such as those utilized by HuggingFace[23, 24], were also considered but not included here due to their specialization and association with proprietary platforms. The evaluation focused on flexible published frameworks that can be adapted, are open, and align closely with the 7 design principles proposed by Heger et al.[1].
Methods
The approach for designing and implementing the Open Datasheets framework involved a multistage process that incorporated feedback from various dataset producers. This approach also included researching existing data documentation frameworks [14, 15, 19, 16, 20] and studying the challenges and opportunities identified in previous studies [14, 1, 15, 16, 17, 18] on data documentation.
Our primary research aim was to identify the areas that needed improvement in practical implementation for open data documentation. To achieve this, we collaborated with data scientists and researchers from the Microsoft AI for Good Research Lab [25], who leverage and publish open datasets for their efforts to solve societal challenges with AI. Through hands-on experience and collaboration, we aimed to fill the gaps in existing frameworks and documentation practices.
The second aim of our approach was to leverage an existing documentation standard to facilitate adoption, avoid reinvention, and focus on bridging the gaps identified in previous studies on responsible AI documentation. We selected a format and standard based on criteria such as simplicity, user usability, and practicality to support non-developers, researchers, and developers.
The third aim was to provide a framework that includes no-code tooling, resources, and guidance that facilitate the understanding of the importance and application of documentation for open datasets.
We engaged with a diverse cohort of open dataset producers from Microsoft, including data scientists, researchers, analysts, and program managers involved in dataset publication and documentation. Their varying perspectives provided valuable insights.
Initial Iteration
The initial evaluation phase occurred during the early development of the Open Datasheets framework web application. We employed the following methodologies:
  • Case Study Analysis: We applied the Open Datasheets framework to three distinct open data publications hosted on GitHub [26, 27, 28]. These projects were selected to represent a range of producers, including non-developer researchers and data scientists.
  • Qualitative Feedback: We collected insights from publishers through observations and conversations. Their feedback helped us understand the importance of guidance for responsible AI documentation and determine what information and features should be included.
  • Usability Testing: We conducted usability testing to identify challenges in the absence of automated features and inline guidance.
Challenges
Identified Publishers experienced difficulties in several areas:
  • Comprehending the value of responsible AI documentation.
  • Deciding on the required information for adequate documentation.
  • Navigating the framework without inline guidance or automation tools.
These observations were consistent with findings from Heger et al.’s research [1], which highlighted similar challenges in documentation tools.
Iterative Improvement
In response to the feedback from the initial iteration, we refined the framework to provide a better balance of features:
  • Enhanced Inline Guidance: We improved the inline help components to instruct users on creating effective responsible AI documentation, providing clear examples and definitions.
  • Automation Features: We introduced automation tools to streamline the documentation process. These tools can infer and prefill certain parts of the datasheets based on the data itself, reducing manual input.
  • User Documentation: We provided comprehensive user documentation, including a step-by-step guide and resources, to help dataset producers navigate the framework and understand best practices for responsible AI documentation.
Design and Implementation Insights
The refined design and implementation were tailored to address the key challenges identified, resulting in a more intuitive and resourceful framework. Automation and guidance components were fine-tuned to ensure dataset producers receive support throughout the documentation process.
The methods employed serve to refine the Open Datasheets framework into a tool that accommodates varied user needs while emphasizing the importance of robust documentation for open datasets. Each step of the process has been scrutinized and enhanced to align with feedback and prior research outcomes.
Design
@@ Methods
The Open Datasheets framework is designed to streamline the documentation of open datasets. It aims to foster the inclusion of crucial information that assists users in comprehending potential biases, privacy concerns, and other elements of responsible AI. This is achieved by striking a balance between automation and thoughtful deliberation in the documentation process [1]. The ultimate objective is to enhance data documentation for open datasets, improve their discoverability, and promote their reliability, fairness, and transparency when used for AI models.
@@
The Open Datasheets format specification is rooted in the Datapackage standard, a product of the Frictionless Data project[29]. This open-source initiative offers tools and standards for data management. The Datapackage specification, accessible on GitHub[30], is a user-friendly JSON based format that acts as a container for describing a set of data. Also, it supports different data types, including complex ones such as tabular and geographic data, allowing users to document the basic information of a dataset in a way that is tailored to their specific needs. This standard has been widely adopted on open data platforms such as https://data.world.
@@
The Open Datasheets framework utilizes the Datapackage format specifications to consistently document basic dataset details. This approach enables users to easily comprehend the information on the datasheets by avoiding specialized terminology and adhering to the standard language and usage of the JSON format. This format is machine-readable and seamlessly integrates into data workflows, thereby enhancing data discovery, access, and usage. The adoption of these specifications fosters compatibility and uniformity in the realm of data documentation.
@@
"name" : "a-unique-human-readable-and-valid-url-identifier",
"title" : "A descriptive title",
@@
Although the Datapackage specification is beneficial for general data documentation, it may not suffice for organizations that need to comply with regulatory standards and ensure responsible AI practices. To cater to this need, the Open Datasheets framework incorporates concepts from “Datasheets for Datasets”[14]. These concepts, such as detailed descriptions of a dataset’s privacy implications to encourage careful reflection, are integrated into a machine-readable format. This additional information is crucial for organizations to make informed decisions about dataset usage and ensure compliance with their policies.
Moreover, the Open Datasheets Framework adheres to the 7 design principles outlined in section 5.2 of Heger et al research[1]. It explicitly connects data documentation with responsible AI through inline guidance based in part on Microsoft’s Aether data documentation template[22]. The framework is practical, integrating with GitHub and providing a user-friendly interface. It adapts to different contexts, allowing customization. Automation is balanced with responsibility, automating foundational metadata extraction while guiding users on responsible AI metadata. It clarifies the target audience to data publishers, focusing on potential users of open datasets. Standardization is achieved through a machine-readable format, and integration with GitHub promotes collaboration. The framework seamlessly integrates into existing tools and workflows, associating documentation with the data publication and development lifecycle on GitHub.
@@
Implementation
To achieve a no-code solution, the framework implements a user-friendly web application on GitHub Pages (Figure 1), a managed service by the GitHub Platform to host static websites on GitHub repositories. This implementation ensures the longevity and support of the framework. The web application features a wizard-style interface that assists in standardizing the documentation according to the framework’s metadata format. It includes metadata parsers for common data file formats on the GitHub platform[31], such as CSV, TSV, JSON, and others. These parsers extract metadata related to the file structure, including field names, types, and sample values. General metadata about the data file, such as filename, encoding, and size, is also extracted. This approach allows data publishers to focus solely on filling in the responsible AI metadata.
@@
The integration with the GitHub Platform, one of the largest open data platforms[31], promotes openness and community. Currently, the web application enables seamless documentation of datasets on GitHub, enhancing discoverability and transparency for datasets hosted on the platform. However, the Open Datasheets framework is not limited to documenting datasets exclusively on GitHub. It provides the flexibility to create documentation for datasets and download the metadata file for publishing on any preferred platform. Moreover, the framework is openly available on GitHub, ensuring easy accessibility for data publishers.
The Open Datasheets framework simplifies the data documentation process for data publishers and promotes the inclusion of responsible AI elements. By automating the extraction of foundational metadata from data files and providing visual guidance for responsible AI considerations, the framework reduces the time and effort required for extensive documentation. This addresses the reluctance of data publishers to write lengthy documentation, as highlighted in Heger et al research [1].
For data users, the Open Datasheets framework offers detailed and machine-readable documentation, enabling informed decisions regarding dataset selection and use. The framework encourages data publishers to document potential biases and privacy implications associated with the dataset. This documentation empowers data users to evaluate the reliability and fairness of the dataset. The comprehensive documentation encouraged by the framework enhances transparency and trustworthiness, allowing data users to assess the dataset’s quality, understand its limitations, and ensure it aligns with their ethical standards and requirements.
@@
The Open Datasheets framework metadata format comprises two main sections: the dataset foundational metadata and the responsible AI metadata. The dataset foundational metadata extends the Datapackage standard and includes essential information such as the dataset’s name, title, licenses, sources, and a comprehensive list of the data resources within the package. On the other hand, the responsible AI metadata builds on the concepts from "Datasheets for datasets" [14] and Microsoft’s Aether Data Documentation Template [22]. It provides information about the data’s origin, processing methods, privacy implications, potential biases, and other relevant aspects. These components are further discussed in the following sections, specifically the Datapackage and Responsible AI sections.
Datapackage
@@
Responsible AI
The Responsible AI module of the Open Datasheets framework focuses on documenting aspects that can impact user privacy, security, and biases that can undermine the trustworthiness of AI systems when used for training AI models. Evaluating datasets is crucial for creating responsible AI systems because the quality of the data directly affects the accuracy and fairness of the resulting system. Biases in the data can lead to biased outcomes, which can have serious consequences for individuals and society[32, 33, 34]. For instance, if an AI system used for hiring decisions is trained on biased data, it may perpetuate discrimination against certain groups.
@@
However, it is important to note that human review will still be necessary to make decisions based on organizational policies. The Open Datasheets framework is not a substitute for human judgment but rather a tool to assist in the evaluation of responsible AI. By combining human expertise with machine-readable metadata, organizations can create more responsible and trustworthy AI systems.
With this objective in mind, the framework encompasses the following responsible AI areas:
  • Privacy
  • Data Access
  • Collection Procedures
  • Processing Procedures
  • Update Procedures
  • Use Cases
Privacy and Use Terms
@@
is a key aspect of privacy assessments. It refers to the protection of personal information from unauthorized access, use, or disclosure[35]. Therefore, having this incorporated into the framework helps the users of the dataset understand the measures that were taken, if any, by the publishers to avoid the disclosure of personal information in their dataset.
@@
in the context of privacy, refers to the level of sensitivity or potential harm associated with certain types of personal information included in a dataset. This includes sensitive attributes such as race, sex, religion, sexual orientation, health information, financial data, and other personally identifiable information (PII)[36, 37]. These sensitive attributes have the potential to cause harm or discrimination if mishandled, accessed, or disclosed without proper consent or safeguards.
@@
Data Access
Including information on how to access a dataset is also essential for privacy and security considerations. One key reason is to clearly communicate whether the data can be accessed anonymously or if registration is required. This information helps potential users understand how to access the data and determine if they meet the necessary criteria to access it.
@@
Another important reason for specifying data access methods in the dataset documentation is to establish clear terms of use. This helps prevent non-qualified access or misuse of the data, preserving its integrity and privacy.
In line with these considerations, the Open Datasheets framework incorporates a module that enables data publishers to clearly specify how people can access the data. Additionally, it enables dataset providers to establish terms of use for the data and maintain data security. By providing this information, data providers can promote transparency, accountability, and responsible data usage.
Collection Procedures
@@
Additionally, documenting data collection procedures enables the replication of the process in the future, which is vital for researchers to validate the results of an AI system trained on that data. By replicating the study and following the documented collection procedures, researchers can confirm that the system is functioning as intended. If the replicated study yields different results, it suggests the need for further investigation to identify potential issues or biases within the original dataset or data collection procedures.
Furthermore, documenting data collection helps data users ensure that the process adheres to ethical and legal standards. The framework takes into consideration the methods, consent forms utilized for data subjects and the contributors of the collection procedures.
@@
describe who performed the data collection, if an individual, an organization or a third-party. This helps data users understand if the data publisher was not the entity that collected the data.
Processing Procedures
@@
Update Procedures
Datasets can be categorized as either static or periodically updated. A static dataset remains unchanged over time, while a periodically updated dataset undergoes regular updates to incorporate new data or revisions. When choosing between these options, researchers consider their specific objectives. If the research requires analyzing historical data or examining trends over a specific period, a static dataset may be suitable. Conversely, if the research involves analyzing real-time or evolving phenomena, a periodically updated dataset would be more appropriate. By documenting the procedures for updating the dataset, data users can gain insights into the timing and nature of the updates. This documentation promotes transparency and builds trust in the data, as users can verify its recency and reliability.
@@
describes the approach for updating the dataset, such as incrementally or a full refresh of the dataset.
@@
Overall, documenting the permissible and impermissible uses of a dataset is a crucial step in ensuring responsible AI systems. It provides guidelines for utilizing the data in a responsible and ethical manner, while also highlighting clearly ill-advised uses.
@@
Conclusion